# Estimating the Influence of Sequentially Correlated Literary Properties in Textual Classification: A Data-Centric Hypothesis-Testing Approach

### v1

@@ 1 Introduction
In computational linguistics, characterizing specific desired properties of a text – such as theme, style, or genre – poses significant challenges due to their abstract and multifaceted nature Rybicki et al. (2016). These parameters encompass a broad range of linguistic elements, making them difficult to quantify and model computationally. Furthermore, these parameters vary widely across authors and texts, posing challenges for developing universal computational models (Serrano et al., 2009). Identifying patterns and connections within these parameters involves deciphering complex semantic and conceptual relationships, which can be context-dependent and subject to individual interpretation. This variability makes it challenging to automate analysis reliably across diverse texts. Additionally, the classification of texts based on these parameters faces similar challenges, as texts often defy traditional categorization and exhibit hybrid or evolving conventions. Additionally, such classification relies on intricate intertextual and extratextual factors, such as cultural context and reader expectations, which are challenging to formalize computationally. Overall, literary works’ abstract and multidimensional nature presents formidable obstacles in computational linguistics, demanding innovative approaches and interdisciplinary collaborations for adequate characterization.
@@
Of particular interest is the discipline of stylometry, which presupposes the possibility of distinguishing between distinct authorial styles (that are independent of thematic content) based on statistical learning analyses Kestemont (2014). Identifying such a stylistic property is helpful in a variety of real-world settings (e.g., forensic linguistics Lambers and Veenman (2009), plagiarism detection (Ainsworth and Juola, 2018), and authorship analysis (Kestemont et al., 2016) – where it has been shown that other textual properties are often misleading in asserting authorship (Stamatatos, 2009, 2018)).
@@
As with many statistical-learning tasks, approaches in stylometry are usually categorized into supervised and unsupervised settings Holmes (1994, 1998); Juola (2006). In the supervised case, several distinct authors and their works are provided (for training), and the task is to attribute texts to one of the authors under consideration, a task known as authorship attribution. Within this context, two distinct assumptions can be made: (1) that the author of each unattributed text is necessarily found within the given set of authors Koppel et al. (2009), and (2) that the composer of an unattributed text need not be among the authors under consideration Koppel et al. (2011). Alternatively, it may be desirable, given a single known author and some prior texts thereof, to verify whether some unknown work was also composed by this author or, instead, by a different one – a task known as authorship verification (e.g., Koppel et al., 2007; Juola, 2006; Kestemont et al., 2016). Other supervised settings include authorship profiling, where the labeling required is not for a distinct author but for a group with some shared profile (e.g., historical era, native language, gender, socio-economic class) Argamon et al. (2009). Furthermore, variants of these tasks include cross-genre settings, where the train set is from a different genre than the test set (e.g., prose-poetry Kestemont et al. (2012)), cross-domain settings (e.g., Facebook-Twitter Markov et al. (2018)), or both (Stamatatos, 2018). Lastly, there exists a variety of different case studies that incorporate different languages, historical eras, text sizes, and more (Schwartz et al., 2013).
@@
As the challenge at hand is identifying a typical style regardless of thematic content, literary features used in stylometry focus on the abundant rather than the scarce (Koppel and Winter, 2014). Unique standing-out words that allude to content are usually ignored, and preferred features are traditionally subsets of the most frequent words and sequences thereof, as these could attest to some authorial style that persists through different works regardless of content (e.g., Holmes, 1994; Argamon et al., 2007; Stamatatos, 2009; Escalante et al., 2011). Even today, when an interpretable analysis is of importance, commonly used features are $n$-gram features of words, characters ($k$-mers), part-of-speech tags (POS), or dependency parsing tags, which are usually applied with a frequency-dependent vectorization such as tf-idf Sammut and Webb (2010).
While some attention has been devoted to a thorough assessment of the utility of different literary features in various settings, most of this attention has focused on variants of the supervised setting of stylometry Koppel et al. (2013); Eder (2015); Juola et al. (2019). In the unsupervised setting, much less has been studied about the effects of different linguistic features and their dependence on textual properties such as genre or language-specificity. Moreover, the focus has been on assessing the contribution of different features and methods in achieving benchmarks. However, to the best of our knowledge, no prior work deals with ascertaining what part of the signal found can verifiably be attributed to different constituents of the literary signal, which can lead to false assertions concerning the results of such analyses.

### v2

@@ 1 Introduction
In computational linguistics, characterizing specific desired properties of a text – such as theme, style, or genre – poses significant challenges due to their abstract and multifaceted nature (Rybicki et al., 2016). These parameters encompass a broad range of linguistic elements, making them difficult to quantify and model computationally. Furthermore, these parameters vary widely across authors and texts, posing challenges for developing universal computational models (Serrano et al., 2009). Identifying patterns and connections within these parameters involves deciphering complex semantic and conceptual relationships, which can be context-dependent and subject to individual interpretation. This variability makes it challenging to automate analysis reliably across diverse texts. Additionally, the classification of texts based on these parameters faces similar challenges, as texts often defy traditional categorization and exhibit hybrid or evolving conventions. Additionally, such classification relies on intricate intertextual and extratextual factors, such as cultural context and reader expectations, which are challenging to formalize computationally. Overall, literary works’ abstract and multidimensional nature presents formidable obstacles in computational linguistics, demanding innovative approaches and interdisciplinary collaborations for adequate characterization.
@@
Of particular interest is the discipline of stylometry, which presupposes the possibility of distinguishing between distinct authorial styles (that are independent of thematic content) based on statistical learning analyses (Kestemont, 2014). Identifying such a stylistic property is helpful in a variety of real-world settings (e.g., forensic linguistics (Lambers and Veenman, 2009), plagiarism detection (Ainsworth and Juola, 2018), and authorship analysis (Kestemont et al., 2016) – where it has been shown that other textual properties are often misleading in asserting authorship (Stamatatos, 2009, 2018)).
@@
As with many statistical-learning tasks, approaches in stylometry are usually categorized into supervised and unsupervised settings (Holmes, 1994, 1998; Juola, 2006). In the supervised case, several distinct authors and their works are provided (for training), and the task is to attribute texts to one of the authors under consideration, a task known as authorship attribution. Within this context, two distinct assumptions can be made: (1) that the author of each unattributed text is necessarily found within the given set of authors (Koppel et al., 2009), and (2) that the composer of an unattributed text need not be among the authors under consideration (Koppel et al., 2011). Alternatively, it may be desirable, given a single known author and some prior texts thereof, to verify whether some unknown work was also composed by this author or, instead, by a different one – a task known as authorship verification (e.g., Koppel et al., 2007; Juola, 2006; Kestemont et al., 2016). Other supervised settings include authorship profiling, where the labeling required is not for a distinct author but for a group with some shared profile (e.g., historical era, native language, gender, socio-economic class) (Argamon et al., 2009). Furthermore, variants of these tasks include cross-genre settings, where the train set is from a different genre than the test set (e.g., prose-poetry (Kestemont et al., 2012)), cross-domain settings (e.g., Facebook-Twitter Markov et al. (2018)), or both (Stamatatos, 2018). Lastly, there exists a variety of different case studies that incorporate different languages, historical eras, text sizes, and more (Schwartz et al., 2013).
@@
As the challenge at hand is identifying a typical style regardless of thematic content, literary features used in stylometry focus on the abundant rather than the scarce (Koppel and Winter, 2014). Unique standing-out words that allude to content are usually ignored, and preferred features are traditionally subsets of the most frequent words and sequences thereof, as these could attest to some authorial style that persists through different works regardless of content (e.g., Holmes, 1994; Argamon et al., 2007; Stamatatos, 2009; Escalante et al., 2011). Even today, when an interpretable analysis is of importance, commonly used features are $n$-gram features of words, characters ($k$-mers), part-of-speech tags (POS), or dependency parsing tags, which are usually applied with a frequency-dependent vectorization such as tf-idf (Sammut and Webb, 2010).
While some attention has been devoted to a thorough assessment of the utility of different literary features in various settings, most of this attention has focused on variants of the supervised setting of stylometry (Koppel et al., 2013; Eder, 2015; Juola et al., 2019). In the unsupervised setting, much less has been studied about the effects of different linguistic features and their dependence on textual properties such as genre or language-specificity. Moreover, the focus has been on assessing the contribution of different features and methods in achieving benchmarks. However, to the best of our knowledge, no prior work deals with ascertaining what part of the signal found can verifiably be attributed to different constituents of the literary signal, which can lead to false assertions concerning the results of such analyses.